1. Excluding stoppages, the speed of a bus is 54 kmph and including stoppages, it is 45 kmph. For how many minutes does the bus stop per hour?


[A] 9


[B] 10


[C] 12


 


[D] 20 

Grade:9

1 Answers

Palash Agarwal
10 years ago

Answer is [b] 10 as,

the time for which the bus stops = the distance reduced/speed per minute

Distance reduced = 54-45=9km

speed per minute= 54/60=9/10

therefore time for which bus stopped = 9/(9/10)=90/9=10minutes

therefore the bus stopped for 10 minutes per hour